OpenStudy (anonymous):

need help please Write this number using standard notation. 4.018 x10^-4 @mathstudent55

OpenStudy (mathstudent55):

10^-4 means move the decimal point 4 places to the left. A negative power of 10 means smaller than 1. 10^-1 = 1/10 = 0.1 10^-2 = 1/100 = 0.01 10^-3 = 1/1000 = 0.001 etc.

OpenStudy (anonymous):

so whats the answer 0.001

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):